19 November 2007

RHIC Startup:

Yesterday, tune feedback and ramp replay go going in Blue, but not yet in Yellow.  37x37 bunches at 90% ramp efficiency.  Yellow tune path was altered to keep away from the 60 Hz line by Steve T.  This was put into the design edit\or for the machine for the tune we want, instead of the model tune.  There was a major confusion about what was the tune of the machine and what to use for tune feedback.  GregM  will address this problem. 

The vacuum at L05 and L07 were fixed, with vacuum better than ever.

Plan for today is to get the experiments magnets on.  Angelika needs some time to set up the collimators.  There are some dead loss monitors that need fixing.  If there is nothing simple in the service building, there will need to be an access to fix the loss monitors.  Not Yet Ready to start giving the owl shift to experiments.  Angelika needs about 4 hours of owl shift tonight to work on collimator set-up.

PHENIX also would like a 15 minute access to move in the second carriage and check for cleanliness for magnet.

STAR would take advantage of a 1 hour access to change some trigger cables to correct timing. 

Jonathan also needs a 1 hour access to tie in the PHENIX Iris Scanner to make it functional. 

RF needs about 1 hour with beam in each ring to do the rebucketing.  Also needs to do phasing at flattop.  Happy to do it at 6 bunches.

There will be an access today from 13:00 - 14:00 to allow for this work.

 


Scheduling Physicist: Mike Sivertz Last updated:
Email:sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020